OpenStudy (anonymous):

what is the absolute value of the number? 3+i

OpenStudy (perl):

sqrt( 3^2 + 1^2 ) = sqrt(10)

OpenStudy (perl):

maybe you should call it modulus, absolute value might not make sense

OpenStudy (anonymous):

interesting... thank you.

OpenStudy (perl):

but, basically you are asking what is the distance of this complex number if we plot it

OpenStudy (perl):

you can plot 3 + i , on a complex plane where the x axis is real , the y axis is imaginary. the complex number 3 + i will have coordinates (3,1).

OpenStudy (perl):

now what is the distance of this point (3,1) from the origin (0,0)

OpenStudy (perl):

In general the complex number a + bi has coordinates (a,b) , here 3 + i = 3 + 1*i so it has coordinates (3,1)